OpenStudy (anonymous):

the last time youbought pizza, 3 pizzas were just enough for 7 people. At tis rate how many whole pizzas should you buy for a party for 33 people?

OpenStudy (mathmagician):

3 pizzas-7 people x pizzas-33 people x=33*3/7=99/7=14.14 pizzas In this situation, I would by 15 pizzas